OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation 64x^3 + 27 = 0 1. 3/4, -3/4 2.-3/4, 3± 3i√3/8 3. no solution 4. 3/4, 12±12i√3/32

OpenStudy (tkhunny):

Have you considered factoring the Difference of Cubes?

OpenStudy (anonymous):

um, I'm not sure how to do that

OpenStudy (michele_laino):

we can rewrite the left side of the equation, like below: \[\Large 64{x^3} + 27 = \left( {4x + 3} \right) \cdot \left( {16{x^2} - 12x + 9} \right)\]

OpenStudy (anonymous):

I see, after that what do I do?

OpenStudy (michele_laino):

so, we can rewrite that equation as below: \[\Large \left( {4x + 3} \right) \cdot \left( {16{x^2} - 12x + 9} \right) = 0\] which is equivalent to these equations: \[\Large \begin{gathered} 4x + 3 = 0 \hfill \\ 16{x^2} - 12x + 9 = 0 \hfill \\ \end{gathered} \] now, the first equation gives the real solution \(x=-3/4\), so the imaginary or complex solutions are given, solving the second equation: \[\Large 16{x^2} - 12x + 9 = 0\]

OpenStudy (anonymous):

So it would be #1?

OpenStudy (michele_laino):

I think no, since option #1 doesn't contain imaginary or complex numbers

OpenStudy (anonymous):

oops duh. So it's 3 or 4

OpenStudy (michele_laino):

x=-3/4 is the real solution. Now we have to solve this equation: \[\Large 16{x^2} - 12x + 9 = 0\] using the standard formula: \[\Large x = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\] where \(a=16,b=-12,c=9\). Please try to solve such quadratic equation

OpenStudy (anonymous):

it's 4